OpenStudy (lgbasallote):

\[\huge \int (u^2 \div 3 \times u + 4u - 2u^3 \times u) du\]

OpenStudy (anonymous):

good luck with that :)

OpenStudy (anonymous):

Can you separate them ??

OpenStudy (lgbasallote):

haha i know right >:))

OpenStudy (lgbasallote):

this is tricky depending on if you follow PEMDAS or VBODMAS

OpenStudy (lgbasallote):

@waterineyes *where* would you separate them?

OpenStudy (anonymous):

From plus and minus sign.. You will get three separate integrals..

OpenStudy (lgbasallote):

how would that look like?

OpenStudy (anonymous):

May be : \[\int\limits (u^2 \div 3 \times u)du + 4 \int\limits u.du - \int\limits (2u^3 \times u)du\]

OpenStudy (lgbasallote):

so basically...the main problem is that \(u^2 \div 3 \times u\) huh

OpenStudy (anonymous):

I think that will result into : \[\frac{u^2}{3} \times u \implies \frac{u^3}{3}\]

OpenStudy (lgbasallote):

hmm you sure it's not \[\frac{u^2}{3u}\]??

OpenStudy (anonymous):

No.. According to PEMDAS, when division and multiplication occur simultaneously then we do the calculations from left to right.. So division is coming first.. Also according to BODMAS, Division comes before Multiplication..

OpenStudy (lgbasallote):

ahh very well. you pass :D exept that last sentence...shouldn't have said that...

OpenStudy (lgbasallote):

except*

OpenStudy (anonymous):

Yeah I know that.. Sorry for that..

OpenStudy (lgbasallote):

lol